A circular surface with a radius of 0.055 m is exp
Answer #1

Radius = r = 0.055 m

magnitude of electric field = E = 1.38 *104 N/C

electric flux = \Phi = 78 N m2 / C

And we know \Phi = E A cos\Theta ,   A = area of circle = \small \Pi r2 = \small \Pi *0.0552 = 9.5033*10-3 m2

\Rightarrow  78 N m2 /C = 1.38*104 * 9.503 *10-3 * cos\Theta

\Rightarrow cos\Theta = 0.5947

\Rightarrow    \Theta = cos-1(0.5947) = 53.50 degrees